# Artificial Intelligence in Image-based Cardiovascular Disease Analysis: A Comprehensive Survey and Future Outlook

### v1

@@ Abstract
AI, Cardiovascular Diseases, Cardiac Image, Deep Learning, Medical Image Analysis, Survey.
@@ 3.1 AI Methods and Pipelines
The U-Net architecture, introduced by [67], is a prominent model in medical image segmentation. Its popularity stems from its flexibility, efficient design, and proven success across various imaging modalities [110]. U-Net comprises two main components: the Encoder and the Decoder. The Encoder path captures semantic and contextual features through downsampling and convolutional blocks, while the Decoder path uses convolutional and upsampling blocks to gradually increase spatial resolution and facilitate pixel-wise classification. U-Net’s defining feature is its skip connections, which bridge each level of the Encoder to the corresponding Decoder level, ensuring the transfer of high-resolution contextual details. This feature promotes the combination of low-level details with high-context information for precise localization. Furthermore, U-Net’s architecture easily integrates with Recurrent Neural Network (RNN) structures. An example of this integration is the use of ConvLSTM [111] by Azad et al. [112], which nonlinearly combines features from both the encoder and the preceding decoder layers. Since its inception, U-Net has established itself as a benchmark in medical image segmentation, inspiring various advanced derivatives that enhance its original design. For a comprehensive review of U-Net variants, see [113]. In addition to U-Net, Generative Adversarial Networks (GANs) also play a significant role in medical image segmentation [114, 115, 116, 117]. More recently, denoising diffusion models, a notable subset of generative models, have recently attracted significant attention in the field of deep learning. They have demonstrated remarkable utility across a wide range of applications, especially in the enhancement of medical image segmentation [118]. These models have demonstrated their effectiveness in generating high-quality segmented images, further expanding the possibilities in medical imaging analysis.
@@ 4.3 Interactions between Heart and Other Organs
Cine-MRI LV Left Ventricle Segmentation and Quantification Spatio-temporal Multi-task Learning [255]
Cine-MRI WH Synthesis of Myocardial Velocity hybrid deep learning digital twin framework [256]
LGE-CMR LV Cardiac Segmentation V-Net [257]
@@ 5.1 Public Datasets
UKB [258]. The UK Biobank (UKB) stands as a monumental biomedical database, offering genetic and health data from half a million UK participants, making it a cornerstone for research on prevalent and life-threatening diseases. It includes cardiac MRI among other imaging data, making it a treasure trove for AI-driven cardiovascular research.
@@ 6 Future Directions
AI Foundation Models for Heart Health. AI is increasingly becoming the primary approach for analyzing CVD through imaging, aiming to reduce costs and the reliance on invasive procedures [201, 197, 194]. AI algorithms have been developed for tasks such as disease classification, risk prediction, treatment planning, and clinical decision-making in cardiovascular imaging [283]. However, most current AI approaches in this field rely on task-specific models [78], which may not fully capture the complexity and diversity of medical data due to their smaller size and task-focused design. The emergence of AI foundation models, particularly prominent in computer vision and natural language processing [284, 285, 286, 287, 288], represents a significant shift. These large-scale models, with millions or billions of parameters, are capable of handling more complex data and identifying a wide range of patterns and relationships, often achieving state-of-the-art results across various tasks due to their scalability and generalization capabilities [289, 290, 291, 292]. In cardiovascular health, foundation models have the potential to be adapted or fine-tuned for specific CVD analysis tasks [293], such as Atrial Fibrillation [20], Ischemic Heart Disease [90], and Coronary Artery Diseases [201]. Adapting these models to the diverse imaging modalities used in CVD analysis presents challenges due to the significant differences from natural images, based on various physics-based properties and energy sources [294]. A potential solution involves training foundation models on both medical and natural images through fine-tuning approaches [295, 296], aiming to provide a robust foundational solution for clinical challenges in CVD. This advancement could significantly improve the effectiveness and efficiency of diagnosing and treating cardiovascular diseases, marking a substantial step forward in the field [297]. Furthermore, developing more efficient neural network architectures is crucial for advancing foundation models. Notably, Mamba [298], a State Space Model, has emerged as a significant innovation for its adeptness at efficiently modeling long-range dependencies in sequential data, attributed to its exceptional memory efficiency and computational speed [299, 300].
@@ Acknowledgment
Xin Wang is supported by SUNY at Albany Start-up Grant. The research of Dr. Hongtu Zhu is partially supported by the National Institute on Aging (NIA) of the National Institute of Health (NIH) under Award Number RF1AG082938.

### v2

@@ Abstract
AI, Cardiac Image, Cardiovascular Diseases, Deep Learning, Medical Image Analysis, Survey.
@@ 3.1 AI Methods and Pipelines
The U-Net architecture, introduced by [67], is a prominent model in medical image segmentation. Its popularity stems from its flexibility, efficient design, and proven success across various imaging modalities [110]. U-Net comprises two main components: the Encoder and the Decoder. The Encoder path captures semantic and contextual features through downsampling and convolutional blocks, while the Decoder path uses convolutional and upsampling blocks to gradually increase spatial resolution and facilitate pixel-wise classification. U-Net’s defining feature is its skip connections, which bridge each level of the Encoder to the corresponding Decoder level, ensuring the transfer of high-resolution contextual details. This feature promotes the combination of low-level details with high-context information for precise localization. Furthermore, U-Net’s architecture easily integrates with Recurrent Neural Network (RNN) structures. An example of this integration is the use of ConvLSTM [111] by Azad et al. [112], which nonlinearly combines features from both the encoder and the preceding decoder layers. Since its inception, U-Net has established itself as a benchmark in medical image segmentation, inspiring various advanced derivatives that enhance its original design. For a comprehensive review of U-Net variants, see [113]. In addition to U-Net, Generative Adversarial Networks (GANs) also play a significant role in medical image segmentation [114, 115, 116, 117]. More recently, denoising diffusion models, a notable subset of generative models, have recently attracted significant attention in the field of deep learning. They have demonstrated remarkable utility across a wide range of applications, especially in the enhancement of medical image segmentation [118]. These models have demonstrated their effectiveness in generating high-quality segmented images, further expanding the possibilities in medical imaging analysis [119].
@@ 4.3 Interactions between Heart and Other Organs
Cine MRI LV Left Ventricle Segmentation and Quantification Spatio-temporal Multi-task Learning [256]
Cine MRI WH Synthesis of Myocardial Velocity hybrid deep learning digital twin framework [257]
LGE CMR LV Cardiac Segmentation V-Net [258]
@@ 5.1 Public Datasets
5.1.1 Dataset without Genetics Data
@@
Stanford Center for Artificial Intelligence in Medicine and Imaging (AIMI) shared dataset [263]. The AIMI dataset offers a repository of clinical imaging data, including echocardiograms and cardiac MRI data, curated for research purposes. These datasets support the development of AI tools for cardiovascular research.
5.1.2 Dataset with Genetics Data
UK Biobank (UKB) [264]. The UKB stands as a monumental biomedical database, offering genetic and health data from half a million UK participants, making it a cornerstone for research on prevalent and life-threatening diseases. It includes cardiac MRI among other imaging data, making it a treasure trove for AI-driven cardiovascular research.
Heart Imaging Genetics Knowledge Portal (Heart-KP) [265]. Heart-KP is a powerful resource for advancing cardiovascular research by integrating heart imaging and genetic data. This portal provides GWAS results for heart imaging traits, including data on left and right ventricle traits, atrium traits, and aorta traits. It supports extensive genetic and phenotypic data analysis for heart-related research. It enables researchers to explore the genetic basis of heart structure and function, contributing to the development of new diagnostic and therapeutic strategies for heart diseases.
American Heart Association (AHA) Data Repositories [263]. The AHA repositories include datasets that provide detailed imaging data for cardiovascular research. These can include MRI, CT scans, echocardiograms, and other types of heart imaging studies. The repositories also offer genetic data, including genome-wide association studies (GWAS), exome sequencing, and other genetic information relevant to cardiovascular health. This data helps in understanding the genetic underpinnings of heart diseases and conditions, which archives and distributes results of studies investigating the interaction of genotype and phenotype, including those related to heart disease.
@@ 6 Future Directions
AI Foundation Models for Heart Health. AI is increasingly becoming the primary approach for analyzing CVD through imaging [286], aiming to reduce costs and the reliance on invasive procedures [202, 198, 195]. AI algorithms have been developed for tasks such as disease classification, risk prediction, treatment planning, and clinical decision-making in cardiovascular imaging [287]. However, most current AI approaches in this field rely on task-specific models [78], which may not fully capture the complexity and diversity of medical data due to their smaller size and task-focused design.
The emergence of AI foundation models, particularly prominent in computer vision and natural language processing [288, 289, 290, 291, 292], represents a significant shift. These large-scale models, with millions or billions of parameters, are capable of handling more complex data and identifying a wide range of patterns and relationships, often achieving state-of-the-art results across various tasks due to their scalability and generalization capabilities [293, 294, 295, 296, 297]. In cardiovascular health, foundation models have the potential to be adapted or fine-tuned for specific CVD analysis tasks [298], such as Atrial Fibrillation [20], Ischemic Heart Disease [90], and Coronary Artery Diseases [202]. Adapting these models to the diverse imaging modalities used in CVD analysis presents challenges due to the significant differences from natural images, based on various physics-based properties and energy sources [299]. A potential solution involves training foundation models on both medical and natural images through fine-tuning approaches [300, 301], aiming to provide a robust foundational solution for clinical challenges in CVD. This advancement could significantly improve the effectiveness and efficiency of diagnosing and treating cardiovascular diseases, marking a substantial step forward in the field [302]. Furthermore, developing more efficient neural network architectures is crucial for advancing foundation models. Notably, Mamba [303], a State Space Model, has emerged as a significant innovation for its adeptness at efficiently modeling long-range dependencies in sequential data, attributed to its exceptional memory efficiency and computational speed [304, 305].
@@ Acknowledgment
Xin Wang is supported by SUNY at Albany Start-up Grant. The research of Dr. Hongtu Zhu is partially supported by the National Institute on Aging (NIA) of the National Institute of Health (NIH) under Award Number RF1AG082938, U01AG079847, and R01AR082684.